On the first day of its fiscal year, Chin Company issued $10,200,000 of five-year, 12% bonds to finance its operations of producing and selling home improvement products. Interest is payable semiannually. The bonds were issued at a market (effective) interest rate of 13%, resulting in Chin Company receiving cash of $9,833,410.

a. Journalize the entries to record the following:

  1. Issuance of the bonds.
  2. First semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)
  3. Second semiannual interest payment. The bond discount amortization, using the straight-line method, is combined with the semiannual interest payment. (Round your answer to the nearest dollar.)

For a compound transaction, if an amount box does not require an entry, leave it blank. Round your answers to the nearest dollar.

1.
2.
3.
Answer #1

The bond is issued at the discount as the amount received on issue of bonds is less than the par value.

Discount = 10,200,000 - 9,833,410 = 366,590

Bond is issued for the 5 years and interest is payable semiannually. The discount is to be amortized over 10 periods using straight line method.

Amortization of discount for every 6 months = 366,590/10 = $36,659

Interest payment every 6 month = 10,200,000*12%*6/12 = 612,000

Interest expense to be recorded every 6 month = 612,000+36,659 = $648,659

Journal entries are as below:

Journal entries in the books of Chin Company Date Particulars Debit ($) Credit ($) Issuance of bond Cash A/C ------------ Dr